you're trying to save to buy a new $170,000 Ferrari. You have $41,000 today that can be invested at your bank. The bank pays 6.0 percent annual interest on its accounts. How long will it be before you have enough to buy the car?

A) 24.16 years B) 24.91 years C) 22.41 years D) 24.41 years E) 24.66 years E)
Answer #1

Amount today*(1+rate)^time=Future Value
Hence,
41000*(1+6%)^t=170000
=>t*log(1.06)=log(170000/41000)
=>t=log(170000/41000)/log(1.06)
=>t=24.40798045

Option D